# Decentralized Stochastic Subgradient Methods for Nonsmooth Nonconvex Optimization

### v1

@@ Abstract
In this paper, we concentrate on decentralized optimization problems with nonconvex and nonsmooth objective functions, especially on the decentralized training of nonsmooth neural networks. We introduce a unified framework, named DSM, to analyze the global convergence of decentralized stochastic subgradient methods. We prove the global convergence of our proposed framework under mild conditions, by establishing that the generated sequence asymptotically approximates the trajectories of its associated differential inclusion. Furthermore, we establish that our proposed framework encompasses a wide range of existing efficient decentralized subgradient methods, including decentralized stochastic subgradient descent (DSGD), DSGD with gradient-tracking technique (DSGD-T), and DSGD with momentum (DSGDm). In addition, we introduce SignSGD employing the sign map to regularize the update directions in DSGDm, and show it is enclosed in our proposed framework. Consequently, our convergence results establish, for the first time, global convergence of these methods when applied to nonsmooth nonconvex objectives. Preliminary numerical experiments demonstrate that our proposed framework yields highly efficient decentralized subgradient methods with convergence guarantees in the training of nonsmooth neural networks.
@@ 1 Introduction
Here $f_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R},i\in[d]:=\{1,\ldots,d\}$ are nonconvex and possibly nonsmooth, and exclusively known to the agent $i$. Moreover, ${\bm{x}}_{i}$ refers to the local variables of the agent $i$. Furthermore, $\mathtt{E}$ denotes the set of edges of the graph $\mathtt{G}$ and represents communication links between agents, in the sense that all the agents communicate solely with their immediate neighbors within the graph $\mathtt{G}$.
Existing works on decentralized optimization
The decentralized optimization problem (DOP) has wide applications in data science and machine learning [12, 42], leading to increasing demands for efficient optimization methods. Motivated by stochastic gradient descent (SGD), the decentralized stochastic gradient descent (DSGD) [30, 27] is one of the most popular methods for solving (DOP). The updating scheme of DSGD can be expressed as
${\bm{x}}_{i,k+1}=\sum_{j=1}^{d}{\bm{W}}(i,j){\bm{x}}_{j,k}-\eta_{k}{\bm{g}}_{i,k},$ (1.1)
where ${\bm{g}}_{i,k}$ refers to the (stochastic) gradient of $f_{i}$ at ${\bm{x}}_{i,k}$. Moreover, ${\bm{W}}\in\mathbb{R}^{d\times d}$ represents the mixing matrix corresponding to the graph $\mathtt{G}=(\mathtt{V},\mathtt{E})$. That is, ${\bm{W}}$ is symmetric, doubly stochastic, and ${\bm{W}}(i,j)=0$ whenever $(i,j)\notin\mathtt{E}$ (see Section 2.2 for details). By denoting the collection of local variables at $k$-th iteration as ${\bm{X}}_{k}\in\mathbb{R}^{n\times d}$, where
@@
and with ${\bm{G}}_{k}=[{\bm{g}}_{1,k},{\bm{g}}_{2,k},\ldots,{\bm{g}}_{d,k}]\in\mathbb{R}^{n\times d}$, we can reformulate the update equation (1.1) into the following compact form,
${\bm{X}}_{k+1}={\bm{X}}_{k}{\bm{W}}-\eta_{k}{\bm{G}}_{k}.$ (1.3)
Based on the update scheme of DSGD, a wide range of efficient decentralized methods have been developed for solving (DOP) under the assumption on differentiability of its objective functions. For example, [49, 19] proposes the DSGD with momentum (DSGDm) by integrating the heavy-ball momentum [32] into the DSGD. Moreover, some recent works introduce auxiliary variables within DSGD to track the exact gradient of the objective function in each agent, such as the DSGD with gradient-tracking technique (DSGD-T) [48], EXTRA [37], etc. To provide theoretical guarantees for these decentralized methods with continuously differentiable objective functions, [27] provides the first theoretical analysis for the convergence rate of DSGD. Moreover, [52, 34, 23, 1, 50] further explore the convergence properties of the DSGD, especially regarding network topology and data heterogeneity. Additionally, [49, 51, 28] analyze the global convergence of the different versions of DSGDm they proposed, while [53, 47, 33, 39] show the convergence properties of DSGD-T for continuously differentiable objectives. Furthermore, some existing works develop decentralized proximal gradient methods for (DOP), under the assumption that $f_{i}$ is composed of a differentiable term $g_{i}$ and a convex regularization term $r_{i}$ for all $i\in[d]$. In this context, the objective function is weakly convex and thus Clarke regular. For example, [52, 38] introduces the proximal decentralized gradient method and establishes its global convergence properties based on the update scheme of DSGD. Moreover, [17] introduces a decentralized proximal method named NEXT based on similar gradient-tracking techniques as DSGD-T. Furthermore, [35, 24] improve the convergence analysis for DSGD-T and prove its convergence for weakly convex objectives. Although achieving nice theoretical results, all the aforementioned works require the Clarke regularity [14] of the objective functions in (DOP). However, as discussed in [8], the nonsmooth activation functions, including ReLU and leaky ReLU, are popular building blocks for model neural networks. More importantly, in the training of neural networks, the widely employed ReLU activation function leads to the nonsmoothness of the loss function. As a result, in the decentralized training of various popular neural networks, the objective functions $\{f_{i}:1\leq i\leq d\}$ in (DOP) are usually nonsmooth and lacking Clarke regularity. When applied to the decentralized training of these nonsmooth neural networks, all the aforementioned works are not capable of providing convergence guarantees for their analyzed decentralized methods. How to establish convergence properties for decentralized stochastic subgradient methods under nonconvex and nonsmooth settings largely remains unexplored.
Stochastic subgradient methods for nonsmooth nonconvex optimization
@@
In the training of nonsmooth neural networks, one major challenge comes from how to compute the (stochastic) subgradients of the loss function. In a wide range of deep learning packages, the subgradients of the loss function are computed by the automatic differentiation (AD) algorithms. These algorithms, utilizing the chain rule, recursively construct generalized subgradients by composing the Jacobians of each block. However, since the chain rule fails for Clarke subdifferential [14], the outputs of the AD algorithm may not be contained in the Clarke subdifferential of the loss function. To characterize the outputs of AD algorithms in differentiating nonsmooth neural networks, [15, 8] introduce the class of path-differentiable functions, which is general enough to cover a wide range of objective functions in real-world applications, including semi-algebraic functions, semi-analytic functions, and functions whose graphs are definable in some o-minimal structures [15]. For path-differentiable functions, [8] introduces the concept of the conservative field as a generalization of their Clarke subdifferential, and shows that the chain rule holds for conservative fields. As a result, although the results of AD algorithms may not be constrained in the Clarke subdifferential of the loss function of a nonsmooth neural network, these results are contained in a specific conservative field of that loss function. Therefore, we employ the concept of the conservative field as a generalization of the Clarke subdifferential to characterize the outputs of AD algorithms throughout our paper. To analyze the convergence properties of stochastic subgradient methods, the ordinary differential equation (ODE) approaches are developed [2, 10, 18, 15], which describe the convergence of the iterates in a subgradient method through the asymptotic behavior of the dynamics of corresponding differential inclusion (i.e., a nonsmooth generalization of ODE). Then based on the concepts of path-differentiable functions and conservative field, a wide range of stochastic subgradient methods are developed for nonsmooth nonconvex optimization problems on a single agent (i.e., (DOP) with $d=1$), while enjoying convergence guarantees from ODE approaches. [15, 7, 8] study the global convergence of SGD method. Moreover, [36, 26, 46] introduce the SGD with heavy-ball momentum and establish its global convergence. In addition, [11] proposes the inertial Newton method (INNA) and shows its global convergence on training nonsmooth neural networks. Furthermore, [31] proves the global convergence of the AdaGrad-Norm, and [45] provides convergence guarantees for a family of ADAM methods. In addition, [21, 22] discuss stochastic subgradient methods for optimization problems over manifolds. Significant progress has been made in the study of stochastic subgradient descent methods within a single agent. This motivates us to explore decentralized stochastic subgradient methods under nonconvex and nonsmooth settings based on the ODE approaches mentioned above. Although [41] establishes the global convergence of DSGD on nonsmooth functions utilizing the ODE approach, their analysis heavily relies on the condition $\partial(\sum_{i\in[d]}f_{i})=\sum_{i\in[d]}\partial f_{i}$. This assumption presumes the validity of the chain rule for $\{f_{i}\}$ in (DOP). As the chain rule does not hold for the Clarke subdifferential of general nonsmooth functions, hence limiting their analysis to Clarke regular functions [14]. How to establish convergence properties for decentralized subgradient methods in minimizing nonsmooth path-differentiable functions remains a compelling open question for further investigations.
Contributions
@@
  • A novel framework for decentralized stochastic subgradient methods We introduce set-valued mappings $\{\mathcal{H}_{i}:i\in[d]\}$, where $\mathcal{H}_{i}:\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ characterizes the update schemes within the $i$-th agent. Then we first introduce the following framework based on the DSGD in (1.3),
@@
    Here for any $i\in[d]$, ${\bm{Z}}_{i,k}\in\mathbb{R}^{m}$ denotes the local copies of the variables at the $i$-th agent, and ${\bm{Z}}_{k}=[{\bm{z}}_{1,k},\ldots,{\bm{z}}_{d,k}]\in\mathbb{R}^{m\times d}$, i.e., the nested local variables. Moreover, ${\bm{H}}_{k}\in\mathbb{R}^{m\times d}$ refers to the noiseless nested update directions of all the agents, in the sense that ${\bm{H}}_{k}{\bm{e}}$ is an approximated evaluation of $\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}({\bm{z}}_{i,k})$ with ${\bm{e}}:=[\frac{1}{d},\ldots,\frac{1}{d}]^{\top}\in\mathbb{R}^{d}$. We show that such a condition is general enough to enclose a wide range of decentralized subgradient methods. Furthermore, $\Xi_{k+1}\in\mathbb{R}^{m\times d}$ refers to the noises, and $\eta_{k}>0$ denotes the step-size. In addition, ${\bm{W}}\in\mathbb{R}^{d\times d}$ refers to the mixing matrix that characterizes the local consensus of all the agents in each iteration. We prove that under mild conditions, $\lim_{k\to+\infty}\left\|{\bm{z}}_{i,k}-{\bm{z}}_{j,k}\right\|=0$ holds for any $i,j\in[d]$, and the sequence $\{{\bm{z}}_{i,k}\}$ asymptotically approximates the trajectories of the differential inclusion
    $\frac{\mathrm{d}{\bm{z}}}{\mathrm{d}t}\in-\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}({\bm{z}})\right).$ (1.4)
@@
    When the differential inclusion (1.4) admits Lyapunov function $\Psi$ and stable set $\mathcal{A}$, we prove that any cluster point of $\{{\bm{Z}}_{i,k}\}$ lies in $\mathcal{A}$ under mild conditions based on the existing ODE approaches.
@@
  • Decentralized stochastic subgradient methods with convergence guarantees We demonstrate that a wide range of decentralized stochastic subgradient methods fit into our proposed framework, including DSGD, DSGDm, and DSGD-T. Furthermore, we also propose a decentralized variant of SignSGD denoted as DSignSGD which can be enclosed by (DSM). More importantly, based on our proposed framework, we provide convergence guarantees for these decentralized subgradient methods. As there is no existing work on the convergence properties of DSGDm, DSGD-T, and DSignSGD in nonsmooth nonconvex optimization, our results fill the gap between their practical implementations and theoretical guarantees. Preliminary numerical experiments show the efficiency of analyzed methods and exhibit the superiority of our proposed method DSignSGD in comparison with DSGD and DSGDm, hence demonstrating the flexibility and great potential of our framework.
Organizations

### v2

# Decentralized Stochastic Subgradient-based Methods for Nonsmooth Nonconvex Optimization

@@ Abstract
In this paper, we concentrate on decentralized optimization problems with nonconvex and nonsmooth objective functions, especially on the decentralized training of nonsmooth neural networks. We introduce a unified framework to analyze the global convergence of decentralized stochastic subgradient-based methods. We prove the global convergence of our proposed framework under mild conditions, by establishing that the generated sequence asymptotically approximates the trajectories of its associated differential inclusion. Furthermore, we establish that our proposed framework covers a wide range of existing efficient decentralized subgradient-based methods, including decentralized stochastic subgradient descent (DSGD), DSGD with gradient-tracking technique (DSGD-T), and DSGD with momentum (DSGD-M). In addition, we introduce the sign map to regularize the update directions in DSGD-M, and show it is enclosed in our proposed framework. Consequently, our convergence results establish, for the first time, global convergence of these methods when applied to nonsmooth nonconvex objectives. Preliminary numerical experiments demonstrate that our proposed framework yields highly efficient decentralized subgradient-based methods with convergence guarantees in the training of nonsmooth neural networks.
@@ 1 Introduction
Here for any $i=1,\ldots,d$, let $f_{i}({\bm{x}}):=\frac{1}{|\mathcal{S}_{i}|}\sum_{l\in\mathcal{S}_{i}}F_{i}({\bm{x}};{\bm{s}}_{l})$, where $\mathcal{S}_{i}$ denotes the local dataset of agent $i$ and $\{{\bm{s}}_{l}:l\in\mathcal{S}_{i}\}$ represents the samples within dataset $\mathcal{S}_{i}$. Each $F_{i}(\cdot,{\bm{s}}_{l}):\mathbb{R}^{n}\rightarrow\mathbb{R},l\in\mathcal{S}_{i}$ is nonconvex and possibly nonsmooth, and exclusively known to the agent $i$. ${\bm{x}}_{i}$ refers to the local variables of the agent $i$. For convenience, we define the sum function $f$ as
$f({\bm{x}}):=\frac{1}{d}\sum_{i=1}^{d}f_{i}({\bm{x}})$ (1.1)
Moreover, $\mathtt{E}$ denotes the set of edges of the graph $\mathtt{G}$ and represents communication links between agents, in the sense that all the agents communicate solely with their immediate neighbors within the graph $\mathtt{G}$.
1.1 Existing works on decentralized stochastic optimization
The decentralized optimization problem (DOP) has wide applications in data science and machine learning [1, 2], leading to increasing demands for efficient optimization methods. Motivated by stochastic gradient descent (SGD), the decentralized stochastic gradient descent (DSGD) [3, 4] is one of the most fundamental methods for solving (DOP). When we assume the differentiability of $\{f_{i}\}$, the update scheme of DSGD can be expressed as
${\bm{x}}_{i,k+1}=\sum_{j=1}^{d}{\bm{W}}(i,j){\bm{x}}_{j,k}-\frac{\eta_{k}}{|\mathcal{B}_{i,k}|}\sum_{l\in\mathcal{B}_{i,k}}\nabla F_{i}({\bm{x}}_{i,k};{\bm{s}}_{l}),$ (1.2)
where the subset $\mathcal{B}_{i,k}\subseteq\mathcal{S}_{i}$ refers to the mini-batch selection of the samples. Therefore, for any $i\in[d]:=\{1,\ldots,d\}$, the term $\frac{1}{|\mathcal{B}_{i,k}|}\sum_{l\in\mathcal{B}_{i,k}}\nabla F_{i}({\bm{x}}_{i,k};{\bm{s}}_{l})$ refers to the stochastic gradient of $f_{i}$ at ${\bm{x}}_{i,k}$, In addition, we define $\xi_{i,k+1}:=\frac{1}{|\mathcal{B}_{i,k}|}\sum_{l\in\mathcal{B}_{i,k}}\nabla F_{i}({\bm{x}}_{i,k};{\bm{s}}_{l})-\nabla f_{i}({\bm{x}}_{i,k})$ can be viewed as the corresponding evaluation noise. Moreover, ${\bm{W}}\in\mathbb{R}^{d\times d}$ represents the mixing matrix corresponding to the graph $\mathtt{G}=(\mathtt{V},\mathtt{E})$. That is, ${\bm{W}}$ is symmetric, doubly stochastic, and ${\bm{W}}(i,j)=0$ whenever $(i,j)\notin\mathtt{E}$ (see Section 2.2 for details).
For better understanding of the existing approaches, we give a unified notation system. We denote the nested local variables ${\bm{X}}_{k}$, evaluation noises $\Xi_{k+1}$, and noiseless gradients at $k$-th iteration ${\bm{G}}_{k}$ as
@@ 1.1 Existing works on decentralized stochastic optimization
$\Xi_{k+1}:=$ $\left[\xi_{1,k+1},\xi_{2,k+1},\ldots,\xi_{d,k+1}\right],$
${\bm{G}}_{k}:=$ $[\nabla f_{i}({\bm{x}}_{1,k}),\ldots,\nabla f_{d}({\bm{x}}_{d,k})],$
Then for any $i\in[d]$, we denote the mapping $\Phi^{\mathrm{GD}}_{i}:=\{\nabla f_{i}\}$ to characterize the local updates in the $i$-th agent in the DSGD method. With this notation, the update scheme of DSGD in (1.2) can be reformulated as follows:
${\bm{X}}_{k+1}={\bm{X}}_{k}{\bm{W}}-\eta_{k}({\bm{G}}_{k}+\Xi_{k+1})$ (1.4)
[…]
Based on DSGD, a wide range of efficient decentralized methods have been developed for solving (DOP) under the assumption of differentiability of its objective functions. For example, [5, 6] propose the DSGD with momentum (DSGD-M) by integrating the heavy-ball momentum [7] into the DSGD. Moreover, some recent works introduce auxiliary variables within DSGD to track the stochastic gradient of the objective function in each agent, such as the DSGD with gradient-tracking technique (DSGD-T) [8, 9], etc.
To provide theoretical guarantees for these decentralized methods with continuously differentiable objective functions, [4] provides the first theoretical analysis for the convergence rate of DSGD. Moreover, [10, 11, 12, 13, 14] further explore the convergence properties of the DSGD, especially regarding network topology and data heterogeneity. Additionally, [5, 15, 16] analyze the global convergence of the different versions of DSGD-M they proposed, while [17, 18, 9, 19] show the convergence properties of DSGD-T for continuously differentiable objectives.
Furthermore, some existing works develop decentralized proximal gradient methods for (DOP), under the assumption that $f_{i}$ is composed of a differentiable term $g_{i}$ and a convex regularization term $r_{i}$ for all $i\in[d]$. In this context, the objective function is weakly convex and thus Clarke regular. For example, [10] introduces Prox-DGD based on the update scheme of DGD, [20] introduces NEXT based on DGD-T and successive convex approximation techniques, while [21] proposes PG-EXTRA derived from EXTRA. The authors all establish the global convergence of their respective algorithms. Moreover, [22, 23, 24, 25] extend the convergence analysis of Prox-DGD and NEXT from the deterministic case to the stochastic case under different assumptions. Additionally, [26] improves the convergence analysis for DSGD-T and proves its convergence for weakly convex objectives.
Although achieving nice theoretical results, all the aforementioned works are capable of handling, at most, non-smooth but Clarke regular [27] objective functions in (DOP). However, as discussed in [28], the nonsmooth activation functions, including ReLU and leaky ReLU, are popular building blocks for modeling neural networks. More importantly, in the training of neural networks, the widely employed ReLU activation function leads to the nonsmoothness of the loss function. As a result, in the decentralized training of various popular neural networks, the objective functions $\{f_{i}:1\leq i\leq d\}$ in (DOP) are usually nonsmooth without Clarke regularity. All these works can not easily extend their convergence theories to the decentralized training of nonsmooth neural networks. How to establish convergence properties for decentralized stochastic subgradient-based methods under nonconvex and nonsmooth settings remains unexplored.
1.2 Stochastic subgradient-based methods for nonsmooth nonconvex optimization
@@
In the training of nonsmooth neural networks, one major challenge comes from computing the (stochastic) subgradients of the loss function. In a wide range of deep learning packages, the subgradients of the loss function are computed by the automatic differentiation (AD) algorithms. Utilizing the chain rule, these algorithms recursively construct generalized subgradients by composing the Jacobians of each block. However, since the chain rule fails for Clarke subdifferential [27], the outputs of the AD algorithms may not be contained in the Clarke subdifferential of the loss function. To characterize the outputs of AD algorithms in differentiating nonsmooth neural networks, [29, 28] introduce the class of path-differentiable functions, which is general enough to cover a wide range of objective functions in real-world applications, including semi-algebraic functions, semi-analytic functions, and functions whose graphs are definable in some o-minimal structures [29]. For path-differentiable functions, [28] introduces the concept of the conservative field as a generalization of their Clarke subdifferential, and shows that the chain rule holds for conservative fields. As a result, although the results of AD algorithms may not be constrained in the Clarke subdifferential of the loss function of a nonsmooth neural network, these results are contained in a specific conservative field of that loss function. Therefore, we employ the concept of the conservative field as a generalization of the Clarke subdifferential to characterize the outputs of AD algorithms throughout our paper.
To analyze the convergence properties of stochastic subgradient-based methods, the ordinary differential equation (ODE) approaches are developed [30, 31, 32, 29], which describe the convergence of the iterates in a subgradient method through the asymptotic behavior of the dynamics of corresponding differential inclusion (i.e., a nonsmooth generalization of ODE). Then, based on the concepts of path-differentiable functions and conservative field, a wide range of stochastic subgradient-based methods are developed for nonsmooth nonconvex optimization problems within a single agent (i.e., (DOP) with $d=1$), while enjoying convergence guarantees from ODE approaches. [29, 33, 28] study the global convergence of SGD method. Moreover, [34, 35, 36] introduce the SGD with heavy-ball momentum and establish its global convergence. In addition, [37] proposes the inertial Newton method (INNA) and shows its global convergence in training nonsmooth neural networks. Furthermore, [38] proves the global convergence of the AdaGrad-Norm, and [39] provides convergence guarantees for a family of ADAM methods. In addition, [40, 41] discuss stochastic subgradient-based methods for optimization problems over manifolds.
Significant progress has been made in studying the convergence properties of stochastic subgradient-based methods within a single agent. This motivates us to explore decentralized stochastic subgradient-based methods under nonconvex and nonsmooth settings based on the above-mentioned ODE approaches. Compared with single-agent settings, establishing the convergence properties of decentralized stochastic subgradient methods is non-trivial, as the update schemes of all the agents are coupled through local communication under decentralized settings. Although [42] establishes the global convergence of DSGD for nonsmooth functions utilizing the ODE approach, their analysis heavily relies on the condition $\partial(\sum_{i\in[d]}f_{i})=\sum_{i\in[d]}\partial f_{i}$. This assumption presumes the validity of the chain rule for $\{f_{i}\}$ in (DOP). However, the chain rule does not hold for the Clarke subdifferential of general nonsmooth functions, hence limiting their analysis to Clarke regular functions [27]. In the presence of limited existing results towards the convergence of decentralized subgradient-based methods in minimizing nonsmooth path-differentiable functions, developing a unified framework for establishing convergence properties for these methods is of great importance in decentralized optimization.
1.3 Contributions
@@
  • A novel framework for decentralized stochastic subgradient-based methods
    Motivated by the update scheme of DSGD (1.4) that characterized by the mappings $\{\Phi^{\mathrm{GD}}_{i}\}$, we first introduce the following unified framework for decentralized stochastic subgradient methods,
@@
    Here, ${\bm{Z}}_{k}=[{\bm{z}}_{1,k},\ldots,{\bm{z}}_{d,k}]\in\mathbb{R}^{m\times d}$ denotes the nested local variables, ${\bm{H}}_{k}\in\mathbb{R}^{m\times d}$ refers to the noiseless nested update directions, $\Xi_{k+1}\in\mathbb{R}^{m\times d}$ refers to the evaluation noises, and $\eta_{k}>0$ denotes the step-size. Moreover, we introduce set-valued mappings $\{\Phi_{i}:\mathbb{R}^{m}\to\mathbb{R}^{m},i\in[d]\}$ to characterize the relationship between ${\bm{Z}}_{k}$ and ${\bm{H}}_{k}$, in the sense that $\frac{1}{d}{\bm{H}}_{k}{\bm{1}_{d}}$ is an approximated evaluation of $\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}({\bm{z}}_{i,k})$. We demonstrate that, with particular choices of the set-valued mappings $\{\Phi_{i}\}$, the framework (DSM) is general enough to enclose a wide range of decentralized subgradient-based methods, including DSGD, DSGD-M, and DSGD-T.
    We prove that under mild conditions, $\lim_{k\rightarrow+\infty}$ $\left\|{\bm{z}}_{i,k}-{\bm{z}}_{j,k}\right\|=0$ holds for any $i,j\in[d]$, and the sequence $\left\{\boldsymbol{z}_{i,k}\right\}$ asymptotically approximates the trajectories of the differential inclusion
    $\frac{\mathrm{d}{\bm{z}}}{\mathrm{d}t}\in-\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}({\bm{z}})\right).$ (1.5)
@@
    When the differential inclusion (1.5) admits Lyapunov function $\psi$ and stable set $\mathcal{A}$, we prove that any cluster point of $\left\{{\bm{z}}_{i,k}\right\}$ lies in $\mathcal{A}$ under mild conditions based on the existing ODE approaches.
  • Convergence guarantees for decentralized stochastic subgradient-based methods
@@
    Based on our proposed framework, we provide convergence guarantees for DSGD, DSGD-M, and DSGD-T. Furthermore, we also propose a decentralized variant of SignSGD [43] denoted as DSignSGD, which can be enclosed by (DSM). As there is no existing work on the convergence properties of DSGD-M, DSGD-T, and DSignSGD in solving nonsmooth nonconvex optimization problems, our results fill the gap between their practical implementations and theoretical guarantees. Preliminary numerical experiments show the efficiency of analyzed methods and exhibit the superiority of our proposed method DSignSGD in comparison with DSGD and DSGD-M, hence demonstrating the flexibility and great potential of our framework.
1.4 Organizations